OpenStudy (anonymous):

What are the zeros of f(x) = x3 + 6x2 + 3x – 10 ?

OpenStudy (anonymous):

do you mean solve for x when f(x) =0?

hero (hero):

Well, at least we know that x = 1 is one of the zeroes, since: f(1) = (1)^3 + 6(1)^2 + 3(1) - 10 f(1) = 1 + 6 + 3 - 10 f(1) = 10 - 10 f(1) = 0 Good luck finding the other two

hero (hero):

If x = 1 is one of the zeroes, then x - 1 is one of the factors. Then all you have to do is long division to find the others.

OpenStudy (anonymous):

is 5 one of the zeros?

hero (hero):

Did you try f(5) ? You should post it on here.

hero (hero):

I'm just going to quit helping you if you don't want to collaborate about this.